OpenStudy (anonymous):

The base of a triangle is 3 centimeters greater than the height. The area is 14cm squared. /find the height and the length if the base.

OpenStudy (anonymous):

\[1/2*(x+3)*(x)=14^{2}\]

OpenStudy (anonymous):

x^2+3x=28 x^2+3x-28=0 solve

OpenStudy (anonymous):

A=bh, and b=h+3. Therefore, 14=h*(h+3), or h^2+3h-14=0. Get yo solve on.

OpenStudy (anonymous):

\[14=\frac{(3+h)*h}{2}\]\[14*2=3h+h^2\]\[28=3h+h^2\]\[h^2+3h-28=0\]Factor:(h+7)(h-4) h= -7 or h= 4 the height can't not negative than the height is : 4 Base = 4+3 Base=7 check: \[\frac{7*4}{2}=\frac{28}{2}=14cm^2\]

OpenStudy (anonymous):

if you like my answer click good answer thank you